# Network-Realised Model Predictive Control Part II: Distributed Constraint Management

### v4

@@ Abstract
A two-layer control architecture is proposed, which promotes scalable implementations for model predictive controllers. The top layer acts as both reference governor for the bottom layer, and as a feedback controller for the regulated network. By employing set-based methods, global theoretical guarantees are obtained by enforcing local constraints upon the network’s variables and upon those of the first layer’s implementation. The proposed technique offers recursive feasibility guarantees as one of its central features, and the expressions of the resulting predictive strategies bear a striking resemblance to classical formulations from model predictive control literature, allowing for flexible and easily customisable implementations.
@@ 1.1 Setting and Context
Model Predictive Control (MPC, see [RMD] and also [BBM]) is one of the most popular control-based frameworks in both the industrial setting and in system-theoretical literature. One of the key features that has generated its current popularity is its affinity for enforcing set-based constraints in the centralised setting [STMC]. On the other hand, as technical applications are becoming increasingly interconnected and widely distributed, the synthesis of distributed control laws has evolved into one of the foci of contemporary research for the control community.
Among the myriad approaches developed for the MPC paradigm (see [DMPCB] for a comprehensive survey on the topic), some of the most promising solutions are those inspired by robust MPC (see [RMPC] for a popular centralised setting), among which we highlight the techniques proposed in [R8A, R8B, R8C, R8D] for the distributed setting. A key aspect of this particular framework is given by the use of an ancillary control law, based upon static state feedback, which operates in tandem with its MPC policy.
Distinct from the latter approach, the procedure formalised in [DMPC1, DMPC2] aims to leverage the powerful design framework from [SLS] in order to tackle the intricacies of the distributed setting. One such intricacy, which arises chiefly in formation control, is the problem of attenuating the propagation of disturbance through said formation (see [SS_rev]). This problem has received a great deal of attention in MPC-oriented literature, with multiple solutions having been proposed in both the distributed setting [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4], and also in the centralised one [plat_MPC_5].
@@ 1.2 Motivation
Despite the valuable contributions made in the previously cited works, certain limitations which are inherent in these approaches serve as the chief motivation for the developments presented in this paper. More precisely, the control laws proposed in [R8A, R8B, R8D] notably require a set of static and stabilising state-feedback strategies, which are constrained to having block-diagonal sparsity structures. Furthermore, the procedure described in [R8C] revolves around intricate set-theoretical notions, such as joint set invariance, while implicitly assuming that the controlled network is bipartite and that the state matrix of its (linear) dynamics satisfies a restrictive spectral condition. Finally, the design framework presented in [DMPC1, DMPC2] tackles both dynamical disturbance decoupling and distributed constraint satisfaction jointly, in a single design phase, whereas the formation control-oriented approaches in [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5] require restrictive or computationally demanding procedures to reach their goals.
@@ 1.3 Scope of Work
In contrast to the techniques from [R8A, R8B, R8C, R8D], the MPC-based control laws discussed in this paper do not focus on network regulation in the presence of constraints, but rather on closed-loop supervision of both the plant and its control laws, which based upon the Network Realisation Function (NRF, see the companion paper [part1], along with [NRF]) representation. In our framework, disturbance rejection and reference tracking are handled predominantly by the distributed (and dynamical) controllers discussed in [part1]. By exploiting the closed-loop properties ensured by these NRF-based subcontrollers [plutonizare, NRF, aug_sparse], the MPC-inspired procedures stated in the sequel are more akin to distributed versions of the so-called reference governors (see [refgov] for a broad discussion on these policies). Indeed, a careful inspection of the survey presented in [arch_rev] reveals that our framework does not fit neatly into any of the architectures addressed therein.
@@ 1.4 Contribution
    coordinating the control authority of our MPC-based policies with the NRF-based controllers from [NRF, aug_sparse], while splitting the MPC’s actuation efforts between the latter and the controlled network;
@@
A further point to emphasise is the fact that all of the above features are achieved in a fully distributed setting, and we point out that the mechanism by which a dynamical controller’s output is reliably constrained (for the purpose of command limitation) has not, to the best of our knowledge, been adequately addressed in system-theoretical literature, even in the centralised setting.
@@ 1.5 Paper Structure
The rest of the paper is structured as follows. Section 2 contains a set of preliminary notions, while Section 3 discusses our two-layer architecture and its overarching objectives, focusing on the MPC layer. Section 4 presents the means of obtaining tractable prediction models for an area-partitioned control scheme, and Section 5 handles the technicalities of constraint management in the distributed setting. Section 6 holds the main theoretical results and implementation procedures, while Section 7 showcases a numerical example based upon the vehicle platooning application discussed in [plutonizare] and Section 8 offers a set of concluding remarks. In addition to this, we include a pair of appendices: one focusing on the key concepts of NRF-based control laws (Appendix A), as discussed in our companion paper [part1], and one which holds the proofs of this paper’s main results (Appendix B).
@@ 2.3 Distributed Networks
Although we consider that the vectors $x$ and $u$ are completely accessible, in terms of measurement and actuation, respectively, we do not assume that manipulating all of these variables is possible from any one location in our physical network. This naturally separates the network into a number of $N\in\mathbb{N}$ distinct areas, with $N>1$.
@@
$S_{\bullet i}:=\left[\scriptsize\begin{array}[]{ccc}e_{\alpha_{\bullet i}+1}&\dots&e_{\alpha_{\bullet i}+n_{\bullet i}}\end{array}\right],\,\forall\,i\in\{1:N\},\vskip-5.69054pt$ (6)
with which we define $x_{i}[k]:=S^{\top}_{{xi}}x[k]$, $x_{c{i}}:=S^{\top}_{{xi}}x_{c}$ and $u_{i}[k]:=S^{\top}_{{ui}}u[k]$. Once the network is partitioned, the design of our control scheme can begin.
@@ 3.1 Global Perspective and Functioning
The aim of this manuscript and of its companion paper [part1] is to formalise the design framework of the two-layer control architecture for the interconnected dynamical system depicted in Figure 1. By identifying the subset of areas $\mathcal{N}_{i}\subseteq\{1:N\}$ which are able to transmit information to the network’s $i^{\text{th}}$ area (and which trivially satisfy $i\in\mathcal{N}_{i}$), for all $i\in\{1:N\}$, the communication graph is clearly designated and the fundamental operating principle of our control scheme is the following:
@@
[…]
@@
As previously mentioned in Remark 2.1, disturbance plays a key role in our chosen framework. Thus, we recall the disturbance vector $d[k]$ from (1a)-(1c) and we denote the measurement noise mentioned in said remark as $\zeta[k]\in\mathbb{R}^{n_{x}}$. Moreover, we also introduce $\beta_{f}[k]\in\mathbb{R}^{n_{u}}$, for the first layer’s communication disturbance, along with $\beta_{s1}[k]\in\mathbb{R}^{n_{x}}$ and $\beta_{s2}[k]\in\mathbb{R}^{n_{u}}$, the communication disturbance associated with the second layer. The latter three signal vectors act as additive disturbance on the command signals depicted in Figure 1, namely $u_{f}[k]:=\tiny\begin{bmatrix}u_{f1}^{\top}[k]&\dots&u_{fN}^{\top}[k]\end{bmatrix}^{\top}$, $u_{s1}[k]:=\tiny\begin{bmatrix}u_{s11}^{\top}[k]&\dots&u_{s1N}^{\top}[k]\end{bmatrix}^{\top}$ and $u_{s2}[k]:=\tiny\begin{bmatrix}u_{s21}^{\top}[k]&\dots&u_{s2N}^{\top}[k]\end{bmatrix}^{\top}$, respectively. All of the exogenous signals mentioned in this paragraph may be combined into the global disturbance vector
@@
with respect to which we now proceed to formally define the objectives of our MPC layer.
@@ 2. II)
    The area-based components of $u_{s1}[k]$ and $u_{s2}[k]$ are computed using receding horizon-based optimisation (see, for example, Chapter 1 in [RMD]), with each area admitting tractable formulations of type (8), located at the top of this page, where:
@@ 4. IV)
As previously indicated in Remark 3.1, the goals stated above show that our second layer is only charged with enforcing constraints in a distributed manner. Thus, all objectives pertaining to reference tracking or robust stabilisation are to be handled by the NRF-based first layer, in order to take full advantage of its remarkable closed-loop guarantees (see, for example, the main results of [plutonizare, NRF, aug_sparse]). In our framework, the MPC layer acts primarily as a safeguard for the NRF-based control laws, further highlighting the particular nature of our approach. As shown by the numerical example in the sequel, we aim to design our second layer so as to avoid acting on the first-layer-based closed-loop system as much as possible, only generating command signals of non-negligible amplitude when constraint violation becomes imminent.
In order to achieve all of the above-stated objectives, however, tractable and area-based models will have to be derived for the first-layer-based closed-loop system.
@@ 4 Obtaining an MPC-Oriented Formulation
This section is dedicated to discussing the derivation and the manipulation of the models which form the equality constraints of the problems stated in (8). Readers familiar with the robust MPC framework (see Chapter 3 in [RMD] for a comprehensive treatment of the topic) will undoubtedly recognise many of the formulations and conventions adopted in the sequel. Nevertheless, the usage of dynamical feedback instead of the classical, static gains in the so-called tube-based MPC framework (see, for example, Section 3.5 in [RMD]) renders our problem distinct from other approaches in system-theoretical literature. Therefore, given this distinct element of novelty, we proceed to lay out our expressions in an explicit manner.
@@ 4.1 Interfacing with the First Layer
We begin by briefly summarising several key concepts from our companion paper [part1]. Our architecture’s first layer is implemented as depicted in Figure 2, located on the next page, in which the distributed control laws are represented by the TFM
@@
We operate under the assumption that the communication infrastructure of our control architecture is endowed with error-detecting/correcting transmission mechanisms. Thus, all communication errors affecting transmitted information are modelled as disturbance which originates at the source of transmission, not at the receiving ends. Consequently, the resulting signal diagram of the first layer’s closed-loop system will be precisely the one shown in Fig. 2, even when all $\mathbf{K}_{\mathbf{D}i}(z)$ are implemented separately, and the closed-loop mappings will have the beneficial properties discussed in [NRF, part1, aug_sparse].
$\mathbf{K}_{\mathbf{D}i}(z)=\left[\scriptsize\begin{array}[]{c|c}A_{wi}-zI_{n_{wi}}&B_{wi}\\ \hline\cr C_{wi}&D_{wi}\end{array}\right]:=\left[\tiny\begin{array}[]{ccc|c}A_{r(\alpha_{ui}+1)}-zI_{n_{r(\alpha_{ui}+1)}}&&&B_{r(\alpha_{ui}+1)}\\ &\ddots&&\vdots\\ &&A_{r(\alpha_{ui}+n_{ui})}-zI_{n_{r(\alpha_{ui}+n_{ui})}}&B_{r(\alpha_{ui}+n_{ui})}\\ \hline\cr C_{r(\alpha_{ui}+1)}&&&O\\ &\ddots&&\vdots\\ &&C_{r(\alpha_{ui}+n_{ui})}&O\\ \end{array}\right].$ (10)
Crucially, these subcontrollers are implemented via the state-space realisations given in (10), located at the top of the next page, whose component matrices (see Section 4.1 in [part1] for more details) are given, $\forall\,\ell\in\mathcal{A}_{ui}$, by
@@
Henceforth, we define the orders $n_{wi}:=\textstyle\sum_{\ell=1}^{n_{ui}}n_{r\ell}$ and $n_{w}:=\textstyle\sum_{i=1}^{N}n_{wi}$. We also denote the state vectors of the implementations from (10) as $w_{i}[k]$ and their initial conditions as $w_{ci}\in\mathbb{R}^{n_{wi}}$. Moreover, we concatenate the latter into $w_{c}:=\tiny\begin{bmatrix}w_{c1}^{\top}&\dots&w_{cN}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$, which are the initial conditions of the first layer’s global implementation.
@@
One of the main features of this control scheme is that it is able to independently compute the control signals of the first layer as follows
@@
for all $i\in\{1:N\}$, while employing only information from each area’s neighbourhood. Algebraically, this reduces to the fact that, for all $j\in\{1:N\}\setminus\,\mathcal{N}_{i}$, the implementations from (10)-(11) satisfy
$\mathrm{col}_{q}(B_{r\ell})=O,\,\forall\,\ell\in\mathcal{A}_{ui},\,q\in\mathcal{A}_{xj}\cup(\{n_{x}\}\oplus\mathcal{A}_{uj}).\vskip-8.53581pt$ (13)
@@
Although we consider the first layer’s control laws to be obtained through an NRF-based framework, as in our companion paper [part1], note that any set of distributed subcontrollers implemented akin to (9)-(13) may be employed for the procedures described in the sequel.
@@
where $\mathbf{F}_{\mathbf{Q}}(z)$ is a strictly proper TFM, which is to say that $\lim_{z\rightarrow\infty}\mathbf{F}_{\mathbf{Q}}(z)=O$. Notably, $\mathbf{F}_{\mathbf{Q}}(z)$ models the closed-loop’s response to exogenous inputs, while $\mathcal{I}_{\mathbf{Q}}[k]$ is a time-varying matrix which governs the closed-loop system’s evolution with respect to its initial conditions.
@@
While the signal vectors $d[k]$ and $\zeta[k]$ in (7) and (implicitly) in (14) possess classical interpretations in system-theoretical literature, the three vectors denoted generically as $\beta_{\bullet}[k]$ in (7) are caused primarily by encoding errors in the employed communication infrastructure. Due to finite-precision representation, the absolute values of the signals in $\beta_{\bullet}[k]$ may be considered to be upper bounded and, unless stated otherwise, the samples making up $\beta_{\bullet}[k]$ are implicitly assumed to be uniformly distributed within their respective bounds. In addition to this, we point out that boundedness (not necessarily in terms of amplitude) also extends to both $d[k]$ and $\zeta[k]$, in most practical applications. Indeed, due to the stabilising properties of NRF-based control, the effect of all these signals upon the closed-loop system can be made bounded in some sense, by an appropriate choice of systemic norm optimisation (see Chapter 4 of [zhou] and Section 5.2 of [part1]). Moreover, the results from [PUB] can be neatly incorporated into our considered approach, if one wishes to delve further into the stochastic setting.
The main result of our companion paper [part1] shows that $\mathbf{F}_{\mathbf{Q}}(z)$ and $\mathcal{I}_{\mathbf{Q}}[k]$ are easily tunable via a freely chosen $\mathbf{Q}$-parameter (hence, the employed notation), thus enabling the complete design of the first layer’s closed-loop response. The full details pertaining to the synthesis and tuning of these closed-loop maps are given in our companion paper [part1], yet, for the benefit of the reader, we have included a summary of the key properties ensured by NRF-based control in Appendix A.

### v5

@@ Abstract
A two-layer control architecture is proposed, which promotes scalable implementations for model predictive controllers. The top layer acts as both a reference governor for the bottom layer and as a feedback controller for the regulated network. By employing set-based methods, global theoretical guarantees are obtained by enforcing local constraints upon the network’s variables and upon those of the first layer’s implementation. The proposed technique offers recursive feasibility guarantees as one of its central features, and the expressions of the resulting predictive strategies bear a striking resemblance to classical formulations from model predictive control literature, allowing for flexible and easily customisable implementations.
@@ 1.1 Setting and Context
Model Predictive Control (MPC, see [19] and also [5]) is one of the most popular control-based frameworks in both the industrial setting and in system-theoretical literature. One of the key features that has generated its current popularity is its affinity for enforcing set-based constraints in the centralised setting [4]. On the other hand, as technical applications become increasingly interconnected and widely distributed, the synthesis of distributed control laws has emerged as a focus of contemporary research in the control community.
Among the myriad approaches developed for the MPC paradigm (see [15] for a comprehensive survey on the topic), some of the most promising solutions are those inspired by robust MPC (see [16] for a popular centralised solution), among which we highlight the techniques proposed in [8, 26, 22, 11] for the distributed setting. A key aspect of this particular framework is given by the use of an ancillary control law, based upon static state feedback, which operates in tandem with its MPC policy.
Distinct from the latter approach, the procedure formalised in [1, 2] leverages the powerful design framework from [29] to address the intricacies of the distributed setting. One such intricacy, which arises chiefly in formation control, is the problem of attenuating the propagation of disturbance through said formation (see [9]). This problem has received a great deal of attention in MPC-oriented literature, with multiple solutions having been proposed in both the distributed setting [31, 28, 6, 10], and also in the centralised one [27].
@@ 1.2 Motivation
Despite the valuable contributions made in the previously cited works, certain limitations inherent in these approaches serve as the chief motivation for the developments presented in this paper. More precisely, the control laws proposed in [8, 26, 11] notably require a set of static and stabilising state-feedback strategies, which are constrained to having block-diagonal sparsity structures. Furthermore, the procedure described in [22] revolves around intricate set-theoretical notions, such as joint set invariance, while implicitly assuming that the controlled network is bipartite and that the state matrix of its (linear) dynamics satisfies a restrictive spectral condition. Finally, the design framework presented in [1, 2] tackles both dynamical disturbance decoupling and distributed constraint satisfaction jointly, in a single design phase, while the formation control-oriented approaches in [31, 28, 6, 10, 27] require restrictive or computationally demanding procedures to reach their goals.
@@ 1.3 Scope of Work
In contrast to the techniques from [8, 26, 22, 11], the MPC-based control laws discussed in this paper do not focus on network regulation in the presence of constraints, but rather on closed-loop supervision of both the plant and its control laws, which based are upon the Network Realisation Function (NRF, see the companion paper [24], along with [21]) representation. In our framework, disturbance rejection and reference tracking are handled predominantly by the distributed (and dynamical) controllers discussed in [24]. By exploiting the closed-loop properties ensured by these NRF-based subcontrollers [20, 21, 25], the MPC-inspired procedures described in the sequel are more akin to distributed versions of the so-called reference governors (see [7] for a broad discussion of these policies). Indeed, a careful inspection of the survey presented in [23] reveals that our framework does not neatly fit into any of the architectures addressed therein.
@@ 1.4 Contribution
    coordinating the control authority of our MPC-based policies with the NRF-based controllers from [21, 25], while splitting the MPC’s actuation efforts between the latter and the controlled network;
@@
A further point to emphasise is the fact that all of the above features are achieved in a fully distributed setting. Moreover, we point out that the mechanism by which a dynamical controller’s output is reliably constrained (for command limitation) has not, to the best of our knowledge, been adequately addressed in system-theoretical literature even in the centralised setting.
@@ 1.5 Paper Structure
The rest of the paper is structured as follows. Section 2 contains a set of preliminary notions, while Section 3 discusses our two-layer architecture and its overarching objectives, focusing on the MPC layer. Section 4 presents the means of obtaining tractable prediction models for an area-partitioned control scheme, and Section 5 handles the technicalities of constraint management in the distributed setting. Section 6 holds the main theoretical results and implementation procedures, while Section 7 showcases a numerical example based upon the vehicle platooning application discussed in [20] and Section 8 offers a set of concluding remarks. In addition to this, we include a pair of appendices: one focusing on the key concepts of NRF-based control laws (Appendix A), as discussed in our companion paper [24], and one which holds the proofs of this paper’s main results (Appendix B).
@@ 2.3 Distributed Networks
Although we consider that the vectors $x$ and $u$ are completely accessible, in terms of measurement and actuation, respectively, we do not assume that manipulating all of these variables is possible from any one location in our physical network. The latter fact naturally separates the network into $N\in\mathbb{N}$ distinct areas, with $N>1$.
@@
$S_{\bullet i}:=\left[\scriptsize\begin{array}[]{ccc}e_{\alpha_{\bullet i}+1}&\dots&e_{\alpha_{\bullet i}+n_{\bullet i}}\end{array}\right]\in\mathbb{R}^{n_{\bullet}\times n_{\bullet i}},\,\forall\,i\in\{1:N\},\vskip 2.84526pt$ (6)
for which $x_{i}[k]:=S^{\top}_{{xi}}x[k]$, $x_{c{i}}:=S^{\top}_{{xi}}x_{c}$, $u_{i}[k]:=S^{\top}_{{ui}}u[k]$.
@@ 3.1 Global Perspective and Functioning
The aim of this manuscript and of its companion paper [24] is to formalise the design framework of the two-layer control architecture for the interconnected dynamical system depicted in Figure 1. By identifying the subset of areas $\mathcal{N}_{i}\subseteq\{1:N\}$ which can transmit information to the network’s $i^{\text{th}}$ area (and which trivially satisfy $i\in\mathcal{N}_{i}$), for all $i\in\{1:N\}$, the communication graph is clearly designated and the fundamental operating principle of our control scheme can be expressed as follows:
@@
$\min\limits_{u_{s1i}[k],\dots,u_{s1i}[k+T_{i}-1],u_{s2i}[k],\dots,u_{s2i}[k+T_{i}+\overline{T}_{i}-1]}\textstyle\sum_{t=1}^{T_{i}+\overline{T}_{i}}g_{t}(\xi_{i}[k+t],u_{s1i}[k+t-1],u_{s2i}[k+t-1]),$ (8)
$\text{subject to}\left\{\begin{aligned} &\xi_{i}[k+t]=A_{si}\xi_{i}[k+t-1]+B_{s1i}u_{s1i}[k+t-1]+B_{s2i}u_{s2i}[k+t-1],\,\forall\,t\in\{1:T_{i}+\overline{T}_{i}\},\\ &\xi_{i}[k+t]\in\Xi_{it},\,\forall\,t\in\{1:T_{i}\},\ \xi_{i}[k]=O,\\ &u_{s1i}[k+t-1]\in{\mathcal{U}}_{s1i},\,u_{s2i}[k+t-1]\in{\mathcal{U}}_{s2i},\,\forall\,t\in\{1:T_{i}\}.\end{aligned}\right$.
@@
As previously mentioned in Remark 2.1, disturbance plays a key role in our chosen framework. Thus, we recall the disturbance vector $d[k]$ from (1a)-(1c) and we denote the measurement noise mentioned in said remark as $\zeta[k]\in\mathbb{R}^{n_{x}}$. Moreover, we also introduce $\beta_{f}[k]\in\mathbb{R}^{n_{u}}$, for the first layer’s communication disturbance, along with $\beta_{s1}[k]\in\mathbb{R}^{n_{x}}$ and $\beta_{s2}[k]\in\mathbb{R}^{n_{u}}$, the communication disturbance associated with the second layer. The latter three signal vectors act as sources of additive disturbance on the command signals depicted in Figure 1, namely $u_{f}[k]:=\tiny\begin{bmatrix}u_{f1}^{\top}[k]&\dots&u_{fN}^{\top}[k]\end{bmatrix}^{\top}$, $u_{s1}[k]:=\tiny\begin{bmatrix}u_{s11}^{\top}[k]&\dots&u_{s1N}^{\top}[k]\end{bmatrix}^{\top}$ and $u_{s2}[k]:=\tiny\begin{bmatrix}u_{s21}^{\top}[k]&\dots&u_{s2N}^{\top}[k]\end{bmatrix}^{\top}$, respectively. All of the exogenous signals mentioned in this paragraph may be combined into the global disturbance vector defined as
@@
with respect to which we now formally define the objectives of our MPC layer.
@@ 2. II)
    The area-based components of $u_{s1}[k]$ and $u_{s2}[k]$ are computed using receding horizon-based optimisation (see, for example, Chapter 1 in [19]), with each area admitting tractable formulations of type (8), located at the top of this page, where:
@@ 4. IV)
As previously noted in Remark 3.1, the goals stated above indicate that our second layer is solely responsible for enforcing constraints in a distributed manner. Thus, all objectives related to reference tracking or robust stabilisation are to be handled by the NRF-based first layer, so as to take full advantage of the latter’s remarkable closed-loop guarantees (see, for example, the results in [20, 21, 25]). In our framework, the MPC layer primarily serves as a safeguard for the NRF-based control laws, further highlighting the specific nature of our approach. As shown via the numerical example in the sequel, we design our second layer with the aim of minimising its impact on the first-layer-based closed-loop system, generating command signals of non-negligible amplitude only when constraint violation becomes imminent.
To achieve all of the above-stated objectives, however, tractable and area-based models will have to be derived for the first-layer-based closed-loop system.
@@ 4 Obtaining an MPC-Oriented Formulation
This section is dedicated to discussing the derivation and the manipulation of the models which form the equality constraints of the problems stated in (8). Readers familiar with the robust MPC framework (see Chapter 3 in [19] for a comprehensive treatment of the topic) will undoubtedly recognise many of the formulations and conventions adopted in the sequel. Nevertheless, the use of dynamical feedback rather than classical static gains in the so-called tube-based MPC framework (see, for example, Section 3.5 in [19]) makes our problem distinct from other approaches in system-theoretical literature. Therefore, given this distinct element of novelty, we proceed to lay out our expressions in an explicit manner.
@@ 4.1 Interfacing with the First Layer
We begin by briefly summarising several key concepts from our companion paper [24]. Our architecture’s first layer is implemented as depicted in Figure 2, located on the next page, in which the distributed control laws are represented by the following TFM:
@@
We assume that the communication infrastructure of our control architecture is equipped with error-detection/correction mechanisms. Thus, all communication errors affecting transmitted information are modelled as additive disturbance signals which originate at the source of transmission, not at the receiving ends. Consequently, the resulting signal diagram of the first layer’s closed-loop system will be precisely the one shown in Fig. 2, even when all $\mathbf{K}_{\mathbf{D}i}(z)$ are implemented separately, and the closed-loop mappings will have the beneficial properties discussed in [21, 24, 25].
$\mathbf{K}_{\mathbf{D}i}(z)=\left[\scriptsize\begin{array}[]{c|c}A_{wi}-zI_{n_{wi}}&B_{wi}\\ \hline\cr C_{wi}&D_{wi}\end{array}\right]:=\left[\tiny\begin{array}[]{ccc|c}A_{r(\alpha_{ui}+1)}-zI_{n_{r(\alpha_{ui}+1)}}&&&B_{r(\alpha_{ui}+1)}\\ &\ddots&&\vdots\\ &&A_{r(\alpha_{ui}+n_{ui})}-zI_{n_{r(\alpha_{ui}+n_{ui})}}&B_{r(\alpha_{ui}+n_{ui})}\\ \hline\cr C_{r(\alpha_{ui}+1)}&&&O\\ &\ddots&&\vdots\\ &&C_{r(\alpha_{ui}+n_{ui})}&O\\ \end{array}\right].\vskip 5.69054pt$ (10)
Crucially, these subcontrollers are implemented via the state-space realisations given in (10), located at the top of the next page, whose component matrices (see Section 4.1 in [24] for more details) are given, $\forall\,\ell\in\mathcal{A}_{ui}$, by
@@
Henceforth, we define the orders $n_{wi}:=\textstyle\sum_{\ell=1+\alpha_{ui}}^{\,n_{ui}+\alpha_{ui}}n_{r\ell}$ and $n_{w}:=\textstyle\sum_{i=1}^{N}n_{wi}$. We also denote the state vectors of the implementations from (10) as $w_{i}[k]$ and their initial conditions as $w_{ci}\in\mathbb{R}^{n_{wi}}$. Moreover, we concatenate the latter into $w_{c}:=\tiny\begin{bmatrix}w_{c1}^{\top}&\dots&w_{cN}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$, which are the initial conditions of the first layer’s global implementation.
@@
One of the main features of this control scheme is the fact that it can independently compute the control signals of the first layer as follows:
@@
for all $i\in\{1:N\}$, while employing only information from each area’s neighbourhood. This reduces to the fact that the implementations from (10)-(11) must satisfy
$B_{r\ell}\,\mathrm{diag}(S_{uj},S_{xj})=O,\,\forall\,\ell\in\mathcal{A}_{ui},\,j\in\{1:N\}\setminus\,\mathcal{N}_{i}.\vskip-8.53581pt$ (13)
@@
Although we consider the first layer’s control laws to be obtained through an NRF-based framework, as in our companion paper [24], note that any set of distributed subcontrollers implemented akin to (9)-(13) may be employed for the procedures described in the sequel.
@@
where $\mathbf{F}_{\mathbf{Q}}(z)$ is a strictly proper TFM, which is to say that $\lim_{|z|\rightarrow\infty}\mathbf{F}_{\mathbf{Q}}(z)=O$. Notably, $\mathbf{F}_{\mathbf{Q}}(z)$ models the closed-loop’s response to exogenous inputs, while $\mathcal{I}_{\mathbf{Q}}[k]$ is a time-varying matrix which governs the closed-loop system’s evolution with respect to its initial conditions.
@@
While the signal vectors $d[k]$ and $\zeta[k]$ in (7) and (implicitly) in (14) possess classical interpretations in system-theoretical literature, the three vectors denoted generically as $\beta_{\bullet}[k]$ in (7) are caused primarily by encoding errors in the employed communication infrastructure. Due to finite-precision representation, the absolute values of the signals in $\beta_{\bullet}[k]$ may be considered to be upper bounded and, unless stated otherwise, the samples making up $\beta_{\bullet}[k]$ are implicitly assumed to be uniformly distributed within their respective bounds. In addition to this, we point out that boundedness (not necessarily in terms of amplitude) also extends to both $d[k]$ and $\zeta[k]$, for most practical applications. Indeed, due to the stabilising properties of NRF-based control, the effect of all these signals upon the closed-loop system can be made bounded in some sense, by an appropriate choice of systemic norm optimisation (see Chapter 4 of [32], and Section 5.2 of [24]). Moreover, the results from [12] can be neatly incorporated into our considered approach if one wishes to delve further into the stochastic setting.
The main result of our companion paper [24] shows that $\mathbf{F}_{\mathbf{Q}}(z)$ and $\mathcal{I}_{\mathbf{Q}}[k]$ are easily tunable via a freely chosen $\mathbf{Q}$-parameter (hence, the employed notation), thus enabling the complete design of the first layer’s closed-loop response. The full details regarding the synthesis and tuning of these closed-loop maps are given in our companion paper [24], yet, for the benefit of the reader, we have included a summary of the key properties ensured by NRF-based control in Appendix A.